in economic theory, a " hurdle rate" is the minimum return that a person requires before they will make an investment. a research report says that annual returns from specific class of common equities are distributed according to a normal distribution with mean of 12 percent and a standard deviation of 18 percent. a stock screener would like to identify a hurdle rate such that only 1 in 20 equities is above that value. where should the hurdle rate be set?
Answer #1

z = 2-1 Za x-12 The the two found expression of the z-score then 8 -12 -=1-645 18 og a-12=1.645 X 18 12-41.61 Thus the hurdle